I am given the B leg and the hypotenuse the B leg being 6 meters and the hypotenuse being (2x-6) meters. It wants me to find the lengths of two legs and the hypotenuse. I've tried it for the past hour now and cant figure it out..
Found 2 solutions by jorel555, Boreal:
Answer by jorel555(1290)

The hypotenuse is (2x-6) meters, and one leg is 6 meters. So the missing leg,x, would be:
(2x-6)^2-6^2=x^2
4x^2-24x+36-36=x^2
3x^2=24x
x^2=8x
x=8
Your triangle is 6x8x10!!!!!!!!!!!!!

Answer by Boreal(15235)

The square of the A leg is the hypotenuse squared minus 36, B squared
That is 4x^2-24x+36-36=4x^2-24x
A^2=4x^2-24x
A= sqrt (4x^2-24x)
This must be greater than 0, 4x^2>24x, x>6,
Therefore, x must be greater than (6);
There are infinite possibilities greater than (6)
if x=7, we have a 6, sqrt(28) and 8 triangle, the squares on both sides=64
if x=10, we have a 6, sqrt (160) and 14 triangle.
If two sides are both x, there are infinite ways they can work to form a right triangle. They still have to fit with the Pythagorean theorem.